OpenStudy (tavia):

7. In a multiple-choice test, each question has four options. Students will get 2 points for each correct answer; lose 1 point for each incorrect answer; and receive no points for unanswered questions. A student does not know the correct answer for one question. Is it to her advantage or disadvantage to guess an answer? Show your calculations for expected value.

OpenStudy (amistre64):

expected value is determined as: gain x probability

OpenStudy (amistre64):

if they answer they expected value is\[2\times\frac{1}{4}-1\times\frac{3}{4}\] if they dont answer they expected value is just 0 i believe

OpenStudy (tavia):

ok could you explain how you got that

OpenStudy (amistre64):

i thought i did ...

OpenStudy (amistre64):

how many options do you get for a question?

OpenStudy (tavia):

4

OpenStudy (amistre64):

good, and how many options are correct?

OpenStudy (tavia):

2

OpenStudy (amistre64):

2 is the value of a correct option; how many correct options are there to a question?

OpenStudy (tavia):

1

OpenStudy (amistre64):

then the probability of making a correct guess is: 1/4 and the expected value is to gain 2 points 2 * 1/4

OpenStudy (amistre64):

how many options are incorrect?

OpenStudy (tavia):

1

OpenStudy (amistre64):

no, if you have 4 options, and only 1 is correct; how many does that leave to be incorrect?

OpenStudy (tavia):

3

OpenStudy (amistre64):

then the probability of guess incorrectly is: 3/4 and the gain for each incorrect options is: -1 the expected value of an incorrect guess is then: -1*3/4

OpenStudy (amistre64):

the expected value of the whole question is the sum of both of these

OpenStudy (amistre64):

\[2\times \frac14-1\times \frac34=-\frac14\] 2-1-1-1 = -1, out of 4 options

OpenStudy (amistre64):

since they "expect" to lose 1 point by guessing, and 0 points by leaving it blank; should they try to guess at it?

OpenStudy (tavia):

no

OpenStudy (amistre64):

so is guessing advantageous? or disadvantageuos?

OpenStudy (tavia):

disadvantageous

OpenStudy (amistre64):

correct

OpenStudy (tavia):

thanks